OpenStudy (anonymous):

can anyone explain to me how to solve this problem? y= x^2-3x-2 y= x-5

OpenStudy (gorv):

we have y=x-5 so put y value in other equation x-5=x^2-3x-2 now solve it ...give it a try

OpenStudy (gorv):

wlcm to OS :)

OpenStudy (anonymous):

thank you(: